Grade 12th passPhysical Chemistry

Two liquid A and B have v.p ratio 1:3 at certain temperature assume A and B are ideal solution and ratio of mole fraction of A and B in vapour phase is 4:3 then mole fraction of B in solution is what

bhawna
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
Total vapour pressure = 1+ 3= 4
Mole fraction of A = 4/7
Mole fraction of B = 3/7
P(A) = X(A) * Ptotal = 4*4/7 = 16/7
P(B) = 3* 4/7 = 12/7
Xb = 12/7 / 16/7 = 12/16 = 3/4